# Multi-Scale Fusion for Object Representation

### v1

@@ 4.2 VAE Guidance
To investigate how MSF fuses multi-scale information into VAE guidance, we adopt the object separability analysis in (Stanic et al., 2024; Lowe et al., 2024). Intuitively, better object separability in VAE guidance contributes to better OCL performance. Given the number of objects in current image or video frame, we conduct clustering on the VAE guidance; Then we can calculate the inter/intra-cluster distances for quantitative results; Similarly, we also visualize the object separability separately of the scale-invariant representation $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ and scale-variant representation $\{\bm{X}_{n}^{\mathrm{sv}}\}$, as well as the representation before inter-scale fusion $\{\bm{X}_{n}^{\mathrm{si}}\}$ and the representation after intra-scale fusion $\{\bm{X}_{n}\}$, i.e., VAE guidance. The visualization scheme follows (Stanic et al., 2024) and (Lowe et al., 2024), and the coloring scheme follows (Caron et al., 2021).
As shown in Fig. 3 and Tab. 4, with our MSF, good object separability can be observed both qualitatively and quantitatively in all scales $n=1...3$; By contrast, the object separability of the basis methods is consistently worse. Generally, as shown in Fig. 3, (i) image pyramid and shared codebook quantization equips $\{\bm{X}_{n}^{\mathrm{si}}\}$ with naive scale invariance; (ii) inter-scale fusion augments $\{\bm{X}_{n}^{\mathrm{si}}\}$ by one another and yields $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$, where different sized objects are all well represented; (iii) non-shared codebook quantization enables $\{\bm{X}_{n}^{\mathrm{si}}\}$ to preserve scale-variance, i.e., more details; (iv) intra-scale fusion combines the scale-invariance of $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ and the scale variance of $\{\bm{X}_{n}^{\mathrm{sv}}\}$ together, and yields our $\{\bm{X}_{n}\}$, which provides better guidance to OCL compared with the naive VAE intermediate representation $\{\bm{X}\}$ of those basis methods.
Figure 3: Qualitative object separability of VAE guidance with our MSF. Take the upper left as an example. Super-pixels of the circled cube in scale 2 is represented in high-quality by VAE encoder, thus can augment the corresponding super-pixels in scale 1, via inter-scale fusion. Accordingly, the scale-invariant representations $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ are obtained. This can combine with the corresponding scale-variant super-pixels $\{\bm{X}_{n}^{\mathrm{sv}}\}$ in the same scale, producing representations $\{\bm{X}_{n}\}$ with good object separability, via intra-scale fusion. Note: quality values are not drawn.
@@ 4.3 Ablation Study
We use SLATE on ClevrTex as an example to evaluate how different designs contribute to our MSF. We use ARI+ARIfg as the metrics because ARI is primarily dominated by background segmentation performance while ARIfg only measures the foreground. We employ naive CNN (Kipf et al., 2022) as the primary encoder to reduce experiment time.
@@ Appendix A Appendix
A.1 Model Architecture
@@
To save computation while maintain strong performance for all methods, we adopt the big-little design used in (Zhao et al., 2024b), which takes small resolution 128 as input while still produce high resolution (128) feature maps. We employ the DINO v1, tiny, with a downsampling rate of 8, as presented in (Caron et al., 2021). We integrate this into a big-little architecture by running DINO in parallel with the previously mentioned naive CNN. The output of DINO is then projected, upsampled, and combined element-wise with the output of the CNN. The architecture is briefly represented as dino()-conv(k3s1p1)-upsample(k8) + naivecnn(). Notably, the weights of DINO are frozen, which retains its pretrained features while allowing the CNN to focus on refining high spatial details.
@@
For diffusion-based OCL, e.g., SlotDiffusion, we follow the architecture described in (Wu et al., 2023). This decoder take VAE representation with noise as input, taking slots, time step as condition, finally produce the reconstruction of the noise being added to the input. The input channel dimension for this module is set to 4.
@@
For the non-grouped codebook design, we adhere to the typical dVAE and VQ-VAE implementations as described in (Singh et al., 2022a) and (Wu et al., 2023), respectively. For the grouped codebook design used in SLATE/STEVE and SlotDiffusion, we take the OGDR as the basis, by applying with our design as mentioned in main body of this paper. Thus except what we describe in the main body, we also adopt annealing residual connection and Gumbel sampling code matching to stabilize training.
Implementation of our MSF described in Sect 3.3 is described in Code 1.
@@
SysBinder is one of the competitors. This module enhances the SlotAttention module in SLATE/STEVE. We follow its original design while keeping the other components consistent with the earlier-described modules. The grouping number is set to four across all experiments that involve this module. We use group number two.
OGDR is another competitors. We use group sizes of two, and leave all the other settings identical to the original paper.
A.2 Dataset Processing

### v2

@@ 4.2 VAE Guidance
To investigate how MSF fuses multi-scale information into VAE guidance, we analyze object separability (Stanic et al., 2024; Lowe et al., 2024). Empirically, better object separability in VAE guidance contributes to better OCL performance. We cluster on the VAE guidance $\bm{X}_{n=1}$ then calculate inter-/intra-cluster distances; We also visualize (Caron et al., 2021) object separability of scale-invariant representations $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ and scale-variant representations $\{\bm{X}_{n}^{\mathrm{sv}}\}$, as well as representations before inter-scale fusion $\{\bm{X}_{n}^{\mathrm{si}}\}$ and representations after intra-scale fusion $\{\bm{X}_{n}\}$.
As shown in Fig. 3 and Tab. 4, our MSF contributes to significant object separability boosts both qualitatively and quantitatively upon basis methods. In contrast, the object separability of basis methods is consistently inferior to ours.
Figure 3: Qualitative object separability in VAE guidance with MSF. Take the upper left as an example. The circled cube $a$ in column $\bm{X}_{n}^{\mathrm{si}}$ row scale1 is vague and noisy, while its corresponding representation $b$ in row scale2 is very clear, thus $b$ can augment $a$, yielding clearer $c$ in column $\dot{\bm{X}}_{n}^{\mathrm{si}}$ row scale1. As $c$ only captures scale-invariant information, $d$ in column $\bm{X}_{n}^{\mathrm{sv}}$ row scale1 supplements scale-variant information, which has too many details and is not object-like. By combining $c$ and $d$, we obtain $e$ in column $\bm{X}_{n}$ row scale1, making this cube much more separable from the background than in column $X$ the corresponding area. Notations are detailed in Sect. 4.2 beginning.
@@ 4.3 Ablation Study
We use SLATE on ClevrTex as an example to evaluate how different designs contribute to our MSF. We use ARI+ARIfg as the metrics because ARI is mostly dominated by background segmentation performance while ARIfg only measures the foreground. We employ naive CNN (Kipf et al., 2022) as the primary encoder to reduce experiment time.
@@ 5 Conclusion
Acknowledgments
We acknowledge the support of the Finnish Center for Artificial Intelligence (FCAI) and the Research Council of Finland through its Flagship program. Additionally, we thank the Research Council of Finland for funding the projects ADEREHA (grant no. 353198), BERMUDA (362407) and PROFI7 (352788). We also appreciate CSC – IT Center for Science, Finland, for granting access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking and hosted by CSC (Finland) in collaboration with the LUMI consortium. Furthermore, we acknowledge the computational resources provided by the Aalto Science-IT project through the Triton cluster.
@@ Appendix A Appendix
A.1 Extended Related Work
Multi-scale & image/feature pyramid. Existing multi-scale methods (i) Either employ an image pyramid (Adelson et al., 1984; Singh & Davis, 2018; Najibi et al., 2019) without information fusion among pyramid levels, wasting useful information among different scales; (ii) Or rely on feature pyramids, namely, FPN and numerous variants (Lin et al., 2017; Tan et al., 2020; Piao et al., 2021), where features of different DNN depths are fused by channel-concat or element-wise-sum, mixing both low- and high-quality representations together. In contrast, our MSF is the first to enable fusion of multiple scales on VAE representations. By leveraging codebook matching (Van Den Oord et al., 2017), we selectively fuse high-quality information among scales, rather than mixing them together. As shown in Tab. 6, our MSF is superior to naive channel-concat or element-wise-sum.
channel-concat element-wise-sum our msf
ARI+ARIfg 92.63 89.18 100.71
Table 6: Effects of different fusions among scales. Model is SLATE; dataset is ClevrTex.
Multi-scale & VAE. Recent works like VAR (Tian et al., 2024) and SPAE (Yu et al., 2024) also build the multi-scale representations upon VAE. But they are different from our MSF as follows. (i) VAR vs MSF: VAR auto-regresses from smaller scales of VAE representation to the larger, but with no information fusion among scales. In contrast, our MSF realizes fusion among different scales on VAE for the first time. (ii) SPAE vs MSF: SPAE relies on multi-modal foundation model CLIP, while our MSF does not. SPAE element-wisely averages multiple scales into one, mixing both low- and high-quality representations together, i.e., no fusion among scales. Our MSF augments any scaled representation with all other scales’ high-quality representations, i.e., fusion among scales.
Relationships among SSLseg, OCL and WMs. As shown in Fig. 4, (1) SSLseg (Self-Supervised Learning segmentation), e.g., HCL (Ge et al., 2023) and VideoCutLER (Wang et al., 2024), focuses on extracting segmentation masks; (2) OCL (Object-Centric Learning), e.g., SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b), represents each object as a feature vector, with segmentation masks as byproducts, kind of overlapping with SSLseg; (3) WMs (World Models), e.g., SlotFormer (Wu et al., 2023a), upon OCL, address downstream tasks like visual reasoning, planning and decision-making.
Figure 4: SSLseg vs OCL vs WMs.
A.2 Extended Experiments
MSF’s effect on different sized objects. We follow COCO’s small/medium/large size splits to evaluate our MSF’s performance. Results of resolution 128x128 are shown below Tab. 7. Our MSF does show better performance on small objects.
mIoUS mIoUM mIoUL overall
SLATE 8.57 26.65 34.57 26.94
+MSF 12.63 28.14 34.83 29.57
Table 7: How MSF peforms on different sized objects. Dataset is COCO instance segmentation.
MSF on OpenImages subset with higher resolution. We use input resolutions 128 and 256 to evaluate on OpenImages subset Ge et al. (2023). Under resolution 128, we use $n$=3 and 4; under resolution 256, we use $n$=3, 4 and 5. Results are shown below Tab. 8. We use ARI+ARIfg as the metrics because ARI mostly reflect how well the background is segmented and ARIfg only measures foreground objects. According to the results, under resolution 128, n=3 is the best choice; under resolution 256, n=4 is the best choice.
resolution 128$\times$128 256$\times$256
$n$ 3 4 3 4 5
ARI+ARIfg 62.07 60.93 64.82 67.58 62.75
Table 8: Effects of input resolution and #scales $n$. Model is SLATE+MSF; dataset is OpenImages subset.
A.3 Model Architecture
@@
To save computation while maintain strong performance for all methods, we adopt the big-little design used in (Zhao et al., 2024), which takes small resolution 128 as input while still produce high resolution (128) feature maps. We employ the DINO v1, tiny, with a downsampling rate of 8, as presented in (Caron et al., 2021). We integrate this into a big-little architecture by running DINO in parallel with the previously mentioned naive CNN. The output of DINO is then projected, upsampled, and combined element-wise with the output of the CNN. The architecture is briefly represented as dino()-conv(k3s1p1)-upsample(k8) + naivecnn(). Notably, the weights of DINO are frozen, which retains its pretrained features while allowing the CNN to focus on refining high spatial details.
@@
For diffusion-based OCL, e.g., SlotDiffusion, we follow the architecture described in (Wu et al., 2023b). This decoder take VAE representation with noise as input, taking slots, time step as condition, finally produce the reconstruction of the noise being added to the input. The input channel dimension for this module is set to 4.
@@
For the non-grouped codebook design, we adhere to the typical dVAE and VQ-VAE implementations as described in (Singh et al., 2022a) and (Wu et al., 2023b), respectively. For the grouped codebook design used in SLATE/STEVE and SlotDiffusion, we take the GDR as the basis, by applying with our design as mentioned in main body of this paper. Thus except what we describe in the main body, we also adopt annealing residual connection and Gumbel sampling code matching to stabilize training.
Implementation of our MSF described in Sect. 3.3 is described in Code 1.
@@
SysBinder is one of the competitors. This module enhances the SlotAttention module in SLATE/STEVE. We follow its original design while keeping the other components consistent with the earlier-described modules. We use group number two.
GDR is another competitors. We use group sizes of two, and leave all the other settings identical to the original paper.
A.4 Dataset Processing